OpenStudy (anonymous):

I need some help understanding clearing fractions. Example: -x+2/5=5-4/3x

OpenStudy (anonymous):

-x+2/5=5-4/3x

OpenStudy (anonymous):

Is the equation a) \(-x + \frac{2}{5} = 5 - \frac{4}{3x}\) b) \(\frac{-x + 2}{5} = 5- \frac{4}{3x}\) c) None of the above.

OpenStudy (anonymous):

a)

OpenStudy (anonymous):

the x is not below the 4

OpenStudy (anonymous):

oh, good. That's even better.

OpenStudy (anonymous):

\[-x + \frac{2}{5} = 5 - \frac{4}{3}x\]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok, so lets start by getting all of our 'like' terms to have the same denominator. To do that, we'll make the -x into a fraction with a 3 on bottom. How can we do that?

OpenStudy (anonymous):

-x/3

OpenStudy (anonymous):

Nearly, but it must maintain equivalence. 3/3 is 1 right?

OpenStudy (anonymous):

-1x/3

OpenStudy (anonymous):

No. Stop guessing and listen. 3/3 = 1. Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, and 1 times any number gives you the same number right? a*1 = a

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so then: \[a = 1*a = \frac{3}{3}*a = \frac{3a}{3} = a\]

OpenStudy (anonymous):

Did you follow that? Sometimes it's hard to read on one line.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, so what is -x as a multiple of 1/3?

OpenStudy (anonymous):

The basic rule of thumb for equivalent fractions is: You can multiply the top and bottom by the same thing and the value doesn't change.

OpenStudy (anonymous):

don't know

OpenStudy (anonymous):

So we have: \[\frac{-x}{1}\] And we want to get a 3 on the bottom, so we multiply the bottom times 3. That means we have to multiply the top by 3 also to keep it having the same value as -x.

OpenStudy (anonymous):

\[\frac{-3x}{3} = -x\] Right?

OpenStudy (anonymous):

-3x/3

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

So lets do the same thing to the 5 on the right side of the equation. We need to get a 5 on bottom, so what would it be.. We have: \[\frac{5}{1}\]

OpenStudy (anonymous):

5/1+2/5

OpenStudy (anonymous):

Ok, we have this (after changing the denominator on -x) \[\frac{-3x}{3} + \frac{2}{5} = 5 -\frac{4x}{3}\]

OpenStudy (anonymous):

Can you use algebra to get all the terms with x on one side, and all the terms without x on the other?

OpenStudy (anonymous):

I appreciate you trying to help me, but I need to see the example I gave solved from the beginning to the end. It isn't working very well showing me a step here and there.

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

One second.

OpenStudy (anonymous):

Starting from \[\frac{-3x}{3} + \frac{2}{5} = 5 - \frac{4x}{3}\] Can you solve for x?

OpenStudy (anonymous):

James are you there?

OpenStudy (anonymous):

-x+2/5=5-4x/3

OpenStudy (anonymous):

Because you have denominators and an = sign, you need to find a common denominator. What is the common denominator using 5 and 3

OpenStudy (anonymous):

15

OpenStudy (anonymous):

So multiply the whole equation by 15.

OpenStudy (anonymous):

Yes.. every term.

OpenStudy (anonymous):

15(-x+2/5)=15(5-4/3x)

OpenStudy (anonymous):

Yep. And distribute.

OpenStudy (anonymous):

(15)(-x) + (15)(2) (15)(5) - (15)(4x) ----- = ------- 5 3

OpenStudy (anonymous):

Cancel out the denominators into the numerators.

OpenStudy (anonymous):

(15)(-x) + (3)(2) = (15)(5) - 5(4x)

OpenStudy (anonymous):

-15x + 6 = 75 - 20x

OpenStudy (anonymous):

Now solve for x in the usual way.

OpenStudy (anonymous):

Blexting, let him do it.

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

James when you are solving and you see x's on both sides of the equation, I like to get rid of the smallest one. Which one is smaller -15x or -20x

OpenStudy (anonymous):

-20x

OpenStudy (anonymous):

so add 20x to both sides

OpenStudy (anonymous):

James, if you have an equation: 5x + 4 = 4 How do you solve for x?

OpenStudy (anonymous):

What do you get James when you +20x to both sides?

OpenStudy (anonymous):

20x-15x+6=75-20x=20X

OpenStudy (anonymous):

THE LAST ONE should be a plus sign

OpenStudy (anonymous):

Good so add like terms.

OpenStudy (anonymous):

-5x+6=75-x

OpenStudy (anonymous):

20x - 15x (20 - 15 = ?)

OpenStudy (anonymous):

-20x + 20x = (-20 + 20 = ?)

OpenStudy (anonymous):

5

OpenStudy (anonymous):

So you have 5x + 6 = 75 (because your -20 + 20 = 0 so you don't have any x's on that side)

OpenStudy (anonymous):

Then what?

OpenStudy (anonymous):

how do you get the x by itself.. do you get rid of the 5 first or the 6 first?

OpenStudy (anonymous):

5x=75-6

OpenStudy (anonymous):

Yes... but what is 75-6 5x = 69

OpenStudy (anonymous):

but isn't the answer suppose to be x=69/5 and how do you get that?

OpenStudy (anonymous):

Yes... divide in you calculator. 13.8

OpenStudy (anonymous):

where are you getting 13.8

OpenStudy (anonymous):

Once again, I appreciate the help, but I need to see the problem solved from beginning to end. It isn't helping trying to get me to figure out something I don't know. I have to see it work.

OpenStudy (anonymous):

would it be 5x-69 x=69÷5 x=69/5

OpenStudy (anonymous):

5x=69 x=69÷5 x=69/5

OpenStudy (anonymous):

Hello, is anybody there?

OpenStudy (anonymous):

I don't appreciate being ignored. Just because I don't understand doesn't mean I should not be helped. If I knew this crap I wouldn't be here in the first place. Thanks for nothing, and I will not recommend this place to anyone.

OpenStudy (anonymous):

James, nobody knows that you entered another comment on this page unless they are sitting here watching your post... You need to go and "ask a question" and someone will see it and respond...